OpenStudy (anonymous):

Find the area of a circle with a diameter of 12 inches. Use 3.14 for pi

OpenStudy (anonymous):

Area Formula = \[\pi * r^2\] where r = radius radius = 1/2 of diameter 12/2 = 6 = radius 6*2 = 36 36 * pi = 113.04 inches^2

OpenStudy (anonymous):

thanks i got one more The volume of a fish tank is 40 cubic feet. If the density is 0.4fish over feet cubed, how many fish are in the tank?

OpenStudy (anonymous):

Sure, no problem. Density = Mass/Volume Volume = 40 cubic feet Density = 0.4 fish per feet Mass = D * V So, 40 * 0.4 = 10

OpenStudy (anonymous):

10 fish that is